# Laugh Now Cry Later: Controlling Time-Varying Emotional States of Flow-Matching-Based Zero-Shot Text-to-Speech

### v1

@@ 1 Introduction
Humans express a wide range of emotions by changing their tone of voice, often accompanied by nonverbal vocalizations (NVs) such as laughter and crying. While current emotional text-to-speech (TTS) systems have made significant advancements [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12], they still lack the ability to generate emotional speech with fine-grained control (e.g. changing the emotion states within a single generated utterance) and with various types of NVs like laughter and crying. In addition, current emotional TTS systems [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12] are typically trained on staged datasets with a limited number of speakers; in extreme cases, some are trained on only one speaker. These TTS models often lack the ability to generate emotional speech for any speaker, a feature critical for applications like speech-to-speech translation systems that need to retain both the emotion and speaker characteristics of the source audio when generating the translated speech.
In this paper, we introduce EmoCtrl-TTS, an emotion-controllable zero-shot TTS system that can generate highly emotional speech with NVs for any speaker. EmoCtrl-TTS generates the speech by mimicking the voice characteristics and emotion presented by an audio sample, referred to as an audio prompt. EmoCtrl-TTS is based on the flow-matching-based zero-shot TTS [13] and utilizes valence and arousal values to mimic the time-varying characteristics of emotions. In addition, it also utilizes laughter embeddings [14], which we find to be effective for generating not only laughter but also other NVs, including crying. Furthermore, by leveraging over 27k hours of highly expressive real-world data through careful data mining, EmoCtrl-TTS achieves significant enhancements in robustness. Comprehensive evaluations demonstrate that EmoCtrl-TTS excels in reproducing the emotions of audio prompts across multiple languages in speech-to-speech translation scenarios. We also show that EmoCtrl-TTS can capture emotion changes, express strong emotions, and generate various types of NVs in zero-shot TTS.
@@ 2.1 Controlling emotion in TTS
The first point is whether the TTS systems can control the fine-grained emotional attributes within one utterance. Such fine-grained control is ideal for many applications, for example, speech-to-speech translation where nuanced emotional changes need to be transferred to the translated speech. However, as shown in the table, most prior works focused on controlling the utterance-level emotion, and only a few works tackled the control of time-varying emotional status. MsEmoTTS [7] used a local emotional strength predictor to estimate syllable-level emotion strength, leveraging it as a condition to control the emotion strength of the generated speech. ELaTE [14] leveraged laughter representation to condition the flow-matching-based zero-shot TTS, and showed superior controllability of laughter generation. However, these works still lack full controllability of the emotional status.
@@ 3.2 NV embeddings
For our proposed framework, it is essential to figure out a suitable embedding that can represent the characteristics of various NVs. In ELaTE [14], an embedding obtained from an off-the-shelf laughter detection model [17, 18]1 was used to control laughter in the zero-shot TTS.
@@ 4.2.1 Objective evaluation metrics
We utilized the following objective metrics. Among them, AutoPCP, EMO SIM, and Aro-Val SIM are closely related to the emotion controllability. [^7] The evaluation codes for EMO SIM and Aro-Val SIM will be released.

### v2

@@ 1 Introduction
Humans express a wide range of emotions by changing their tone of voice, often accompanied by nonverbal vocalizations (NVs) such as laughter and crying. While current emotional text-to-speech (TTS) systems have made significant advancements [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12], they still lack the ability to generate emotional speech with fine-grained control (e.g. changing the emotion states within a single generated utterance) and with various types of NVs like laughter and crying. In addition, current emotional TTS systems [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12] are typically trained on staged datasets with a limited number of speakers; in extreme cases, some are trained on only one speaker. These TTS models often lack the ability to generate emotional speech for any speaker, a feature critical for applications like speech-to-speech translation, that needs to retain both the emotion and speaker characteristics of the source audio when generating the translated speech.
In this paper, we propose EmoCtrl-TTS, an emotion-controllable zero-shot TTS system that can generate highly emotional speech with NVs for any speaker. EmoCtrl-TTS generates the speech by mimicking the voice characteristics and emotion presented by an audio sample, referred to as an audio prompt. EmoCtrl-TTS is based on the flow-matching-based zero-shot TTS [13] and utilizes valence and arousal values to mimic the time-varying characteristics of emotions. In addition, it also utilizes laughter embeddings [14], which we find to be effective for generating not only laughter but also other NVs, including crying. Furthermore, by leveraging over 27k hours of highly expressive real-world data through careful data mining, EmoCtrl-TTS achieves significant enhancements in robustness. Comprehensive evaluations demonstrate that EmoCtrl-TTS excels in reproducing the emotions of audio prompts across multiple languages in speech-to-speech translation scenarios. We also show that EmoCtrl-TTS can capture emotion changes, express strong emotions, and generate various types of NVs in zero-shot TTS. Our contributions are summarized as follows: (1) we propose a framework integrating flow-matching-based zero-shot TTS with NV and emotion embeddings; and (2) we conduct comprehensive experiments to evaluate the emotion-controllable zero-shot TTS, demonstrating the superiority of the proposed method.
@@ 2.1 Controlling emotion in TTS
The first point is whether TTS systems can control the fine-grained emotional attributes within one utterance. Such fine-grained control is ideal for many applications, for example, speech-to-speech translation where nuanced emotional changes need to be transferred to the translated speech. However, as in Table 1, most prior works aimed to control the utterance-level emotion, and only a few works tackled the control of time-varying emotional status. MsEmoTTS [7] used a local emotional strength predictor to estimate syllable-level emotion strength, leveraging it as a condition to control the emotion strength of generated speech. ELaTE [14] leveraged laughter representation to condition the flow-matching-based zero-shot TTS, and showed superior controllability of laughter generation. However, they still lack full controllability of the emotional status.
@@ 3.2 NV embeddings
For our proposed framework, it is essential to figure out a suitable embedding that can represent the characteristics of various NVs. In ELaTE [14], an embedding obtained from an off-the-shelf laughter detection model [17, 18]1 was used to control laughter in zero-shot TTS.
@@ 4.2.1 Objective evaluation metrics
We utilized the following objective metrics. Among them, AutoPCP, EMO SIM, and Aro-Val SIM are closely related to the emotion controllability. [^7] https://github.com/hbwu-ntu/EmoCtrlTTS-Eval